OpenStudy (anonymous):

Find an equation of the line passing through the point (6,5) and perpendicular to the line whose equation is 2y + 3x = 6.

OpenStudy (anonymous):

slope of 2y + 3x = 6 is -3/2 if its perpendicular, that means the slope is 2/3. y = 2/3 x + b Plug in the coordinates to solve for b. Then you got your equation.